OpenStudy (helpz-q):

Hey guys please help me. it's not crucial but i was scrolling the web when i came across this. i could not figure out the answer so here we are:)

OpenStudy (helpz-q):

This is the question of the web:)

OpenStudy (welshfella):

3^4 = 3*3*3*3 3^2 = 3*3

OpenStudy (welshfella):

or by the rules of exponetns 3^4 / 3^2 = 3^(4-2)

OpenStudy (opticalmango81):

THE 4/2 IS THE EXPONENT

OpenStudy (opticalmango81):

THATS ALL I GOT FOE YA LOL

OpenStudy (helpz-q):

Hmm